A train starts from rest with a uniform acceleration of 10 cm sec-2.it's velocity after one minute will be

Grade:12th pass

1 Answers

Vikas TU
3 years ago
Dear student 
Let the acceleration be  a
Initial velocity of body   u=0
v = u+at 
v = 0+ a
1 min = 60 scond 
so, v = 0 + 10*60 
v = 600 cm/sec 
Hope this will help 
Good Luck 
Cheers